OpenStudy (anonymous):

A street sign that is 9 feet tall casts a shadow that is 24 feet long. A nearby tree casts a shadow that is 56 feet long. Use indirect measurement to find the height of the tree.

OpenStudy (nurali):

9/x=24/56 24x=9(56) x=9(56)/24 x=21

OpenStudy (nurali):

Welcome to Openstudy.